OpenStudy (anonymous):

Having trouble understand this concept. Can someone help? Is the following always, sometimes, or never true? 6 + 8x – 9 = 11x + 14 – 3x (1 point) always true sometimes true never true

OpenStudy (anonymous):

collect terms on the right and on the left. if they are identical, then it is always true they are not identical in this case however

OpenStudy (anonymous):

when you collect terms on the left you have \(8x-3\) because \(6-9=-3\) when you collect terms on the right you have \(8x+14\) because \(11x-3x=8x\)

OpenStudy (anonymous):

so now you are looking at \[8x-3=8x+14\] and then the question is, since they are not identical, is it not "always true" so the question is, is it 'sometimes true" or "never true"? you see you have the same number of \(x\)'s on both sides, namely 8 of them. whatever \(x\) is on the left, it must be the same number on the right, so the \(8x\) on the left is identical to the \(8x\) on the right but the numbers are different therefore it is never true, because you cannot subtract 3 from a number and get the same answer if you add 14 to the same number!

OpenStudy (anonymous):

if you had \(8x-3=7x+14\) then it would be "sometimes true" because you have a different number of \(x\)'s on the left and the right